4. A closed irregular surface encloses a net charge of 1.50 x 10 C. What is the net electric flux through the surface?


1
Expert's answer
2022-03-31T11:49:05-0400

The Gauss's law says

"\phi=\oint EdA=\frac{q}{\epsilon_0}"

"\phi=\frac{1.5*10^{10}}{8.85*10^{-12}}=1.7*10^{21}\:\rm N\cdot m^2/C"